Determine if these lines are perpendicular, parallel or neither. How do I solve this problem?
y=x+7
y=-x+2

Answer by Alan3354(69443)

Determine if these lines are perpendicular, parallel or neither. How do I solve this problem?
y=x+7
y=-x+2
---------------
Parallel lines have the same slope. The slopes of perpendicular lines are negative inverses.
Both these equations are in the slope-intercept form, y = mx+b (m = slope)
The slope of the 1st is 1.
The slope of the 2nd is -1, so they're perpendicular. (1 = -1/1)